Question 3 (1 point) According to a survey conducted by Deloitte in 2017, 0.4532 of U.S. smartphone owners have made an effort to limit their phone use in the past. In a sample of 90 randomly selected U.S. smartphone owners, what is the probability that between 37 and 39 (inclusively) will have attempted to limit their cell phone use in the past? 1) 0.7883 2) 0.1404 3) 0.2117 4) 0.0278 5) 0.6376
Answer #1

p = 0.4532

n=90

Normal approximation to binomail

X = n(np , npq) = N(90 * 0.4532 , 90 *0.4532 *(1-0.4532))

= N( 40.788 , 4.722592 ^2)

Z = (X - mean)/sd = (X - 40.788)/4.722592

P(37 <= X < = 39)

=P(36.5 < X< 39.5)    {continuity correction }

= 0.2106

option 3)0.2116 is most close

hence option 3) is correct

Please rate